When he runs against the wind, Ricky can run 16km in 2 h. With the wind, he can run 5/4 the distance in 2 h. Find Ricky's running rate and the rate of the wind. Assume that both are constant.
Answer by joLeeBee08(3)

we can first simplify the given:
Since he runs 16 km in 2hrs his rate if he runs against the wind is 8 kph.
Against the wind, he runs 5/4 of 16 km which is 20 in 2 hrs so his rate is 10 kph. we should remember that the two rates supposed to be equal but was disturbed by the wind so we need to cancel the rate of the wind so we let x be the wind's rate, and the equation should be:
8+x=10-x >>> x=1
So we found out that the wind has 1kph rate. To find Ricky's rate we must add this to his rate against the wind: 8+1=9
so the answers:
wind's rate: 1kph
Ricky's rate: 9kph